OpenStudy (anonymous):

Which of the following values in the set below will make the equation 2x + 1 = 3 true? (Only input the number.) {0, 1, 2, 3, 4}

OpenStudy (anonymous):

@StudyGurl14 @Mehek14 @kevindragonlord

OpenStudy (anonymous):

@mathmath333

Mehek (mehek14):

answer choices?

OpenStudy (anonymous):

none @Mehek14

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 2x + 1 = 3 \hspace{.33em}\\~\\ \end{align}}\) first substract \(1\) from both sides

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (studygurl14):

2x + 1 =3 solve for x. Do what @mathmath333 said.

OpenStudy (anonymous):

HUH? @StudyGurl14 @MATH

OpenStudy (anonymous):

@mathmath333 HUH?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 2x + 1 -1= 3-1 \hspace{.33em}\\~\\ \end{align}}\) like this i meant ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 2x = 2 \hspace{.33em}\\~\\ \end{align}}\) solve for \(x\) now

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} 2x=2 \hspace{.33em}\\~\\~\\~\\~\\ x=\dfrac{2}{2}=\huge 1 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

THANK UUUUUUUUUUU @mathmath333